CONCENTRATION DEPENDENCE OF THE PHOTOCHEMISTRY OF MATRIX-ISOLATED NI-TROMETHANE

ANGELA M. JARAMILLO, Department of Molecular Biology, Princeton University, Princeton, NJ 08544-1014; and \underline{C. A. BAUMANN}, Department of Chemistry, The University of Scranton, Scranton, PA 18510-4626.

The concentration dependence of the photochemistry of matrix-isolated nitromethane and nitromethane-d_{6} was studied. At higher concentrations, reactions involving fragments from adjacent parent molecules lead to photoproducts not observed in gas phase or dilute matrix studies. In particular, the isomeric forms of N_{2}O_{4} lead to N_{2}O and a number of complexes, most notably CO·NO^{+} and CO·NO_{3}^{−}. This effect is less pronounced when the deuterated compound is irradiated. Density functional theory calculations were used in the characterization of the observed photoproducts and the energetics of the pathways by which they were formed.